### v1

# CNIMA: A Universal Evaluation Framework and Automated Approach for Assessing Second Language Dialogues

Abstract
We develop CNIMA (Chinese Non-Native Interactivity Measurement and Automation), a Chinese-as-a-second-language labelled dataset with 10K dialogues. We annotate CNIMA using an evaluation framework — originally introduced for English-as-a-second-language dialogues — that assesses micro-level features (e.g. backchannels) and macro-level interactivity labels (e.g. topic management) and test the framework’s transferability from English to Chinese. We found the framework robust across languages and revealed universal and language-specific relationships between micro-level and macro-level features. Next, we propose an approach to automate the evaluation and find strong performance, creating a new tool for automated second language assessment. Our system can be adapted to other languages easily as it uses large language models and as such does not require large-scale annotated training data.
CNIMA: A Universal Evaluation Framework and
Automated Approach for Assessing Second Language Dialogues
Rena Gao, Jingxuan Wu, Carsten Roever, Xuetong Wu, Jing Wu, Long Lv, Jey Han Lau The University of Melbourne, Parkville, 3052, Australia The Chinese University of Hong Kong, Shatin, NT, Hong Kong, China Central South University, Changsha, 410017, China Tiangong University, Tianjin, 300387, China {rena.gao,carsten}@unimelb.edu.au, {wfyitf,jeyhan.lau}@gmail.com wujingxuan@cuhk.edu.cn, wujing@tiangong.edu.cn, lvlong@csu.edu.en
1 Introduction
In the context of second language (SL) assessment, speaking has always been considered an essential ability for SL speakers Smith et al. (2022); Allwood (2008); Huang et al. (2020), but prior studies have predominantly focused on written language proficiency Paiva et al. (2022). Given the lack of datasets that capture the unique linguistic features of SL speakers in dialogues (especially in open-domain conversations), this leaves conversational interaction assessment under-explored and ultimately contribute to a limited understanding of conversational fluency and interactivity for SL speakers.
An exception is Gao et al. (2024), who propose a two-level framework for assessing the interactivity ability of English-as-a-second-language (ESL) speakers in open-domain conversations. They introduce micro-level word/utterance features (e.g. backchannels) and macro-level interactivity labels (e.g. topic management) and found that the micro-level features are highly predictive of the macro-level labels. Gao et al. (2024) reveal their intricate relationships, showing that certain micro-level features are highly correlated with a particular macro-level feature, e.g. frequent use of appropriate collocations and varied sentence structures (micro-level features) significantly correlate with topic management and tone appropriateness (macro-level labels). However, this research is conducted only in the context of ESL, raising questions about the transferability of the evaluation framework to other languages besides English. Furthermore, they did not propose an automated pipeline for assessing SL dialogue quality, as their predictive models rely on human-annotated micro-level features.
Our work aims to address these shortcomings by (1) developing and testing a Chinese-as-a-second-language (CSL) dialogue dataset using the evaluation framework of Gao et al. (2024); (2) introducing a fully automated pipeline to assess interactivity for CSL conversations. We experiment with different machine learning models, from classical machine learning models (e.g. logistic regression) to large language models (LLMs; e.g. GPT-4o OpenAI (2024)), to automate the prediction of micro/macro-level features and overall dialogue quality scores. Our best system shows strong predictive performance, paving the way for a new tool for assessing SL conversations [^1] The dataset and related codes can be accessed at https://github.com/RenaGao/CSL2024. To summarise:
  • We release CNIMA (Chinese Non-Native Interactivity Measurement and Automation), an annotated CSL dialogue dataset with 10K dialogues, based on the framework of Gao et al. (2024) originally designed for ESL.
  • We test the cross-lingual transferability of the evaluation framework for CSL and find that it is robust across languages, and we further reveal universal and language-specific relationships between the micro-level and macro-level features.
  • We introduce a fully automated approach that predicts micro-level features, macro-level labels and an overall quality score given an input dialogue by exploring a spectrum of models from classical machine learning to LLMs. Our best system demonstrates strong performance, creating a new automatic tool for second language assessment. Importantly, our LLM-based system can be adapted to other languages without requiring annotated data.
2 Related work
2.1 Assessment of SL Spoken Conversation
Mainstream second language assessments in current industrial practice have mainly focused on grammatical accuracy, pronunciation standardization and vocabulary richness; for example, in TOEFL iBT, PTE Academic and Cambridge IELTS test Xu (2018); Paiva et al. (2022); Xu et al. (2021). However, few speaking assessments emphasised the importance of interaction Khabbazbashi et al. (2021) and aspects such as how speakers manage the topics in communication Shaxobiddin (2024), how speakers perform social roles from speaking Chen et al. (2023), and how speakers start and end a talk in an acceptable manner Yap and Sahoo (2024). There are, however, some exceptions. For example, Dai (2022) develop a test rubric for Chinese SL speakers. More recently, Gao and Wang (2024) introduce an interactivity scoring framework inspired by the IELTS speaking assessment. However, these studies only provide a theoretical framework for evaluation and lack an automated pipeline for large-scale assessments.
2.2 Automated Scoring System in SL
Automated scoring systems can improve the efficiency of processing and scoring dialogues compared to manual assessment methods, allowing for real-time feedback Evanini et al. (2017). Some major second language assessment organisations have employed automated speaking scoring, including PTE Pearson Jones and Liu (2023), Duolingo English Test Burstein et al. (2021), and TOEFL Gong (2023), but these automated systems fail to capture the interactive nature of SL conversations. That is, they struggle to offer detailed analyses and insights on the common errors and language usage patterns and where SL speakers struggle the most. This motivates a more explainable framework for SL assessment that can provide better feedback.
The main challenge is in designing an effective evaluation framework. Capturing the nuances of spoken language, such as interaction, attitudes, and cooperation in conversation, is complex to achieve Gao and Wang (2024). Automated scoring systems must handle diverse patterns, which can vary widely among different non-native SL speakers. Recognizing and assessing interactive features like turn-taking, interruptions, and response appropriateness further complicates the process. Ultimately, the dynamic nature of conversations makes it challenging for automated systems to accurately evaluate interactive speaking, necessitating sophisticated algorithms and continuous refinement Cumbal (2024); Engwall et al. (2022).
3 Evaluation Framework
We adapt the dialogue evaluation framework of Gao et al. Gao et al. (2024) for assessing SL dialogues, which considers micro-level features and macro-level interactivity labels. One addition we have made is to introduce an overall dialogue quality score. Figure 1 shows an example of CSL dialogue annotation with the evaluation framework.
The framework has 17 micro-level features, and it can be further broken into token-level (such as ‘Reference Word’: she and ‘Backchannel’: hmm) and utterance-level features (such as ‘Formulaic response’: How’s going). The micro-level features are annotated as spans (i.e. annotators mark text spans corresponding to a feature). For macro-level interactivity labels, there are four: Topic Management, Tone Choice Appropriateness, Conversation Opening and Conversation Closing, and they are annotated as (dialogue) labels. Briefly, topic management refers to the speaker’s ability to control and navigate the flow of topics; tone choice the suitability of the tone; and conversation opening/closing the naturalness in the initial exchange and conclusion of the discussion. A more elaborate definition can be found in Appendix Table 13. For each macro-level interactivity label, the score ranges from 1 to 5 (categorical), and higher scores indicate a more natural and authentic interactivity quality. Note that for tone choice appropriateness, higher scores denote a casual tone, and lower scores indicate a formal tone. [^2] The rationale of assigning higher scores to casual tone is that in our experiments, the conversations are designed to be informal discussions, and appropriate use of casual tone signifies a more natural communicative interaction.
In addition to micro- and macro-level features, we introduce an overall score for each dialogue to measure second language speakers’ interaction abilities in open-domain conversation. Like the macro-level labels, it is scored from 1 to 5 (categorical). The overall score is designed to capture the holistic quality of a conversation, integrating elements like contextualisation, responsiveness, and communicative purposes across the whole conversation, and a high score reflects the speaker’s ability to engage in fluid and meaningful interactions. The full description of each score is detailed in Appendix Table 12.
Figure 1: An example of a CSL dialogue annotated with the micro-level features, macro-level interactivity labels and overall dialogue quality score.
4 CNIMA Development
4.1 CSL Dialogue Collection
We first extend the CSL dialogue dataset developed by Wu and Roever (2021), which has approximately 8,000 dialogues. We followed a similar process, where we recruited 20 learners of Chinese Mandarin at four different proficiency levels. [^3] Proficiency level can be categorized into the following levels: beginner, lower, intermediate, and high Wu and Roever (2021). Each participant was assigned to a dyadic group with a partner of a similar proficiency level. Each pair of participants did one elicited conversation task, in which they discussed an assigned topic, and two role-play activities (three tasks in total; dialogue collection instructions can be found in Appendix LABEL:subapp:SI). [^4] All participants have signed consent approved by an ethics board to agree to audio recording. After collecting the conversations, we recruited in-house workers to segment the conversations based on the discussion topics. [^5] The workers are undergraduate linguistics major students, and they are native Chinese speakers. Including the original 8,000 dialogues from Wu and Roever (2021), our extended CSL dataset has 10,908 dialogues in total; table 1 presents some statistics.
Statistic Count
#dialogues 10,908
#turns (average) 6
#turns (max) 13
#tokens /w token-level features 170,852
#tokens /w utterance-level features 94,516
Table 1: CNIMA Statistics.
Measure Micro-level Macro-level Overall
Features Labels Score
$\alpha$ 0.66 0.67 0.61
$r$ 0.65 0.68 0.62
Table 2: Inter-annotator agreement for micro-level, macro-level features and overall scores.
4.2 Macro/Micro-level Feature Annotation
Given the dialogues, we annotate them for micro-level features, macro-level interactivity labels and overall quality scores based on the evaluation framework introduced in Section 3. To this end, we have recruited twelve postgraduate students who are native Chinese speakers. The authors of this paper first trained the annotators and the annotator training manual can be found in Appendix  LABEL:subapp:MN. During the training, annotators will see one dialogue as an example to understand the requirements and learn how to use the annotation platform (Appendix A.1). After training, each annotator was assigned 950 dialogues, and each dialogue was annotated by two annotators. [^6] The assignment is created in a way such that we have the same pair of annotators for every batch of 40 dialogues. The annotation was conducted over two months, where in the first few weeks we (first author) checked for initial agreement, discussed feedback, and fixed any annotation errors (e.g. missing values in the annotation results) before proceeding with the full annotation. The annotation process was guided by an annotation guide (Appendix LABEL:subapp:MN), which provided definitions and examples for each micro-level feature, macro-level interactivity label and the overall dialogue quality score. For the overall score, in addition to providing the score, the annotators are also asked to write down their reasons for justifying their label.
As explained in Section 3, the micro-level features are annotated as spans, and the macro-level labels and overall quality scores are document (dialogue) labels. To aggregate the span annotations by the two annotators for the micro-level features, for each dialogue and each micro-level feature, we iterate through each turn and select the shorter (longer) span between the two annotators if it is a token-level (utterance-level) feature. [^7] We do this because token-level features (e.g., ‘Reference word’ (She, her, he) or ‘Tense Choice’ (is doing, done, did) tends to be very short and limited to 1 or 2 words and so the shorter spans are likely to be more accurate Greer (2023). The same reasoning applies to utterance-level spans. For the macro-level interactivity labels, in dialogues where we have disagreement between the two annotators, we use the majority label from their larger unsegmented conversation as the ground truth label. [^8] Recall that these dialogues are segmented from a larger, longer conversation. The majority label is the most frequent label when we pool together all the labels from the segmented, smaller dialogues that belong to the same original conversation. Our rationale for deferring to this majority label is that these are cases where it is difficult to determine the correct dialogue level label using the smaller segmented dialogue (this happens more with topic management than other interactivity labels), and so it is best that we look at all the labels across all the segmented dialogues collection that belong to the same conversation. For the overall dialogue quality scores, for each dialogue with disagreement, we (two authors of this paper) manually assess the justification provided by the annotators to determine the ground truth. [^9] Note that when resolving the disagreement, we only consider the justification without looking at their macro-level interactivity labels. This is so we create an unbiased ground truth for the overall score that is independent of the macro-level labels.
To measure annotation quality, for the micro-level features, we measure agreement between the annotators at the token level for each micro-level feature, i.e., we compute agreement statistics based on the presence or absence of the feature as marked by the annotators for each word token. [^10] In other words, the unit of analysis here is a word token, and the output is a binary value for each annotator indicating whether it has been marked for the feature. We calculate Pearson correlation coefficient $r$ Cohen et al. (2009) and Krippendorff’s $\alpha$ Krippendorff (2018) summarise the results in Table 2. The agreement is above 0.6 over all the features/labels, indicating a good consensus between annotators under this framework.
Figure 2: Pipeline for automated scoring of the CSL dialogue on three steps
5 Automation Pipeline
We now propose automating the prediction of the micro-level features (step 1), macro-level interactivity labels (step 2), and overall score (step 3); see Figure 2 for an illustration. We experiment with classical machine learning models (Logistic Regression (LR), Random Forest (RF), and Naïve Bayes (NB)), fine-tuned Chinese BERT Cui et al. (2021), and GPT-4o OpenAI (2024) with prompts to automate these steps. Note that these predictions are done in sequence, where the output from the previous step is used as input for the next step (e.g. when predicting the macro-level labels, we use the predicted micro-level features as input). We partition CNIMA into train, development and test sets (ratio $=$ 7:1:2) for these experiments.
Step 1
This step predicts the spans of the 17 micro-level features given an input dialogue, and we experiment with BERT and GPT-4o here. For BERT, we fine-tune 17 span classifiers (one for each feature) to classify the presence or absence of a feature for each word in the dialogue. For GPT-4o, we do one-shot prompting (i.e. 1 dialogue with the expected output as a demonstration) with instructions to generate spans for the token-level and utterance-level features separately. [^11] In other words, we have 2 prompts for each dialogue.
Step 2
This step predicts the macro-level interactivity labels for topic management, tone choice appropriateness, and conversation opening and closing. Here, we experiment with classical models (LR, RF and NB), BERT and GPT-4o, and for each of them, we train four (one for each interactivity label) 5-class classifiers (each interactivity label has 5 classes). For the classical models, we follow Gao et al. (2024) and convert the micro-level feature spans into normalised counts and use them as input features to train LR/RF/NB to predict the interactivity labels. For BERT and GPT-4o, the input is the dialogue concatenated with the micro-level feature spans. BERT is fine-tuned as a document classifier, and GPT-4o is one-shot prompted with instructions.
Step 3
This step predicts the overall dialogue quality score. As this is a document classification like step 2, we follow largely the same process for building LR, RF, NB, BERT and GPT-4o. Note that the classical models (LR, RF, NB) use only the normalised macro-level interactivity score as input features (in other words, they have only 4 input features), while BERT and GPT-4o uses a concatenation of the dialogue, micro-level feature spans, and macro-level interactivity scores as input.
For more details on configurations and prompts, see Appendix  A.5 for BERT and  A.6 for GPT-4o.
6 Experiments
We first replicate the ESL experiments of Gao et al. (2024) for CSL and compare their findings in Section 6.1. We then assess the performance of our proposed automated approach for SL assessment in Section 6.2.
Models Topic Tone Opening Closing
LR 0.829 0.859 0.821 0.810
RF 0.831 0.816 0.858 0.852
NB 0.846 0.783 0.811 0.846
Table 3: F1 performance of predicting the macro-level features using human-annotated micro-level features.
Language LR RF NB
ESL Code Switching Code Switching Feedback in Next Turn*
Reference Word* Feedback in Next Turn* Formulaic Responses
Feedback in Next Turn* Question-based responses Reference Word*
Formulaic Responses Non-factive Verb Negotiation of Meaning
Tense Choice Reference Word* Tense Choice
CSL Feedback in the Next Turn Subordinate Clauses* Negotiation of Meaning
Noun & Verb Collocation Routinized Resources Noun & Verb Collocation
Tense Choice Reference Word* Routinized Resources
Reference Word* Code-Switching Subordinate Clauses*
Subordinate Clauses* Feedback in Next Turn Reference Word*
Table 4: High impact common micro-level features over the three classifiers for predicting macro-level features with overlapping features in two/three classifiers by Bold/asterisk. ESL results are reproduced from Gao et al. (2024).
6.1 Transferability of the Evaluation Framework from ESL to CSL
Based on the human-annotated micro-level features, Gao et al. (2024) convert them into normalised counts and use them as input features to train an LR, NB and RF classifier to predict macro-level interactivity labels (i.e. topic management, tone appropriateness, conversation opening and closing). They found: (1) strong prediction performance; (2) high-impact micro-level features that are common across all interactivity labels (by interpreting feature importance given by the trained classifiers); and (3) high-impact micro-level features that are specifically predictive for an interactivity label. We follow Gao et al. (2024) to compute the common and specific features for (2) and (3); details in Appendix A.10.
We replicate their experiments here, using our annotated CSL data (CNIMA), to see whether their findings transfer across languages.
For (1), we present the F1 performance of predicting macro-level interactivity labels for CNIMA in Table 3. We see a strong performance, where F1 is over 0.8 in most models over the 4 interactivity labels. These results echo the ESL results in Gao et al. (2024), [^12] Gao et al. (2024) found marginally higher performance for conversation opening/closing (F1 scores are over 0.9) and lower performance for topic management and tone choice appropriateness (F1 scores are a little below 0.8). providing evidence that the evaluation framework is robust across languages.
For (2), we present the results in Table 4. We found that micro-level features such as ‘Feedback in Next Turn’ and ‘Reference Word’ are high-impact features for both ESL and CSL — this underlines their fundamental role in impacting dialogue interactive dynamics, and it is language-universal. Interestingly, however, we also found some differences. For example, ‘Noun & Verb Collocation’ and ‘Routinized Resources’ are strong features only for CSL, and this might be because, in Chinese, these fixed terms of expressions are often used to show social closeness in open-domain conversations Roever and Dai (2021). ‘Code Switching’ and ‘Tense Choice’, on the other hand, are two strong features only for ESL. This is intuitive, as Chinese has no tenses while English tenses correlate with the social expression in communications Lam (2018). For the full comparison results, see Appendix Table 14.
For (3), Table 5 and 6 present the CSL and ESL results, respectively. We found that for topic management, ‘Negotiation of Meaning’ and ‘Question-Based Responses’ are high-impact micro-level features for both ESL and CSL, demonstrating that these are language-universal features important to drive the flow of topics in conversation. For tone appropriateness, we generally see less commonality (exceptions: ‘Feedback in Next Turn’, and ‘Routinized Resources’ which appear in both languages), suggesting that languages tend to use different features for managing tones Zilio et al. (2017). For conversation opening, we see some similarities (e.g. ‘Question-Based Responses’) and also divergences (e.g. ‘Subordinate Clauses’ and ‘Adj./Adv. Expressing’ for ESL and ‘Epistemic Copulas’ and ‘Epistemic Modals’ for CSL). One explanation is that English tends to use adjectives and adverbs to extend a topic (e.g. Totally, I think…), while Chinese prefers modals and copulas {CJK}UTF8gbsn “应该吧,或许呢” (translation: possibly yeah) to control topic change Alduais et al. (2022). For conversation closing, we see a more similar trend, where ‘Collaborative Finishes’ and ‘Backchannels’ are strong features across both ESL and CSL. Despite some language-specific variations, the strategies of ending a conversation are largely similar in human communication Lam (2021).
Topic Tone Opening Closing
Logistic Regression
Negotiation of Meaning Feedback in Next Turn Epistemic Copulas Backchannels
Epistemic Copulas Collaborative finishes Formulaic Responses Collaborative Finishes*
Collaborative Finishes Routinized Resources Question-Based Responses* Epistemic Copulas
Question-Based Responses Formulaic Responses Epistemic Modals Subordinate Clauses
Backchannels* Question-Based Responses Collaborative finishes Formulaic Responses
Naïve Bayes
Non-factive Verb Phrase Epistemic copulas Code-switching Code-switching
Collaborative Finishes Collaborative finishes Feedback in Next Turn Epistemic Modals
Code-switching Impersonal subject + non-factive verb Epistemic Modals Epistemic Copulas
Backchannels* Backchannels Epistemic Copulas Collaborative Finishes*
Epistemic Copulas Question-Based Responses Question-Based Responses* Question-Based Responses
Random Forest
Noun & verb collocation Tense choice Feedback in Next Turn Backchannels
Negotiation of Meaning Backchannels Noun & Verb Collocation Subordinate Clauses
Backchannels* Negotiation of Meaning Collaborative Finishes Formulaic Responses
Question-Based Responses Feedback in Next Turn Question-Based Responses* Collaborative Finishes*
Collaborative Finishes Routinized Resources Formulaic Responses Question-based Responses
Table 5: CSL: High impact interactivity-specific micro-level features. For each interactivity label, bold/asterisk indicates overlapping features in two/three classifiers.
Topic Tone Opening Closing
Logistic Regression
Negotiation of Meaning* Routinized Resources* Epistemic Modals Backchannels*
Subordinate Clauses* Adj./Adv. Expressing Formulaic Responses Adj./Adv. Expressing
Noun&Verb Collocation Feedback in Next Turn* Question-Based Responses* Formulaic Responses
Question-Based Responses Formulaic Responses* Subordinate Clauses* Collaborative Finishes*
Subordinate clauses* Reference Word Adj./Adv. Expressing* Epistemic Copulas
Naïve Bayes
Non-factive Verb Phrase Routinized Resources* Adj./Adv. Expressing* Adj./Adv. Expressing
Question-Based Responses Feedback in Next Turn* Routinized Resources Epistemic Modals
Adj./Adv. Expressing Epistemic Copulas Subordinate Clauses* Backchannels*
Negotiation of Meaning* Question-Based Responses  Epistemic Copulas Collaborative Finishes*
Subordinate clauses* Subordinate Clauses* Question-Based Responses* Question-Based Responses
Random Forest
Negotiation of Meaning* Epistemic Copulas Feedback in Next Turn Feedback in Next Turn
Formulaic Responses Backchannels Subordinate Clauses* Subordinate clauses
Subordinate Clauses* Feedback in Next Turn* Adj./Adv. Expressing* Collaborative Finishes*
Epistemic Copulas Negotiation of Meaning Question-Based Responses* Formulaic Responses
Question-Based Responses Routinized Resources* Formulaic Responses Backchannels*
Table 6: ESL: High impact interactivity-specific micro-level features (reproduced from Gao et al. (2024)). For each interactivity label, bold/asterisk indicates overlapping features in two/three classifiers.
6.2 Evaluation of Automated Pipeline
We now evaluate our automated 3-step approach to predicting the overall dialogue quality score and present the results in Table 7. In terms of model names, each component denotes the model used in a particular step, e.g. “BERT+LR+LR” means we use BERT for step 1 and LR for steps 2 and 3. Also, “GPT4” refers to GPT-4o. For brevity, we only include LR results for the classical models as they all have similar performances. In addition to our 3-step approach, we also include 2 baselines that predict the overall dialogue quality score directly based on the input dialogue: (1) fine-tuned BERT (“BERT (One-step)”); and (2) one-shot GPT-4o with instructions (“GPT4 (One-step)”). We also include a variation where the first step uses human-annotated micro-level features (e.g. “Human+LR+LR”) to understand how much performance degrades when substituting them with predicted features.
Interestingly, the baselines (“BERT (One-step)” and “GPT4 (One-step)”) perform quite poorly, achieving F1 scores of 0.379 and 0.585, respectively. This indicates using the raw dialogue directly for predicting the overall quality is difficult (most studies in NLP, however, follow this setup for dialogue evaluation Finch et al. (2023); Zhao et al. (2022); Yang et al. (2024)). For the variation where we use human-annotated micro-level features (“Human+LR+LR”, “Human+BERT+BERT”, and “Human+GPT4+GPT4”), we see that BERT is generally the best model and LR the worst, which is no surprise, given that BERT is pre-trained. GPT4, however, is not far from BERT, even though it is not fine-tuned. Overall, the performance is encouraging, and we see that the best models achieve over 0.80 F1.
When we look at the fully automated pipeline (bottom 4 rows in Table 7), “‘BERT+BERT+BERT” and “‘GPT4+GPT4+GPT4” perform very strongly (0.807 and 0.791), demonstrating that we have a fully automated system that can reliably assess the quality score of a dialogue. We also notice an interesting trend: When we use either BERT or GPT4 for predicting the micro-level features, LR (i.e. “BERT+LR+LR” and “GPT4+LR+LR”) performs very poorly for predicting the overall score, even though the span prediction performance of BERT and GPT4 for the micro-level features is no+t poor (we will revisit this in Section 6.3). This suggests that the classical machine learning models are less tolerant of noise. [^13] Note that LR does not use the dialogue as input, so any noise in the intermediate features (micro-level or macro-level) will have a bigger impact.
Taking all these results together, we show that when it comes to assessing SL dialogue quality, it is important to take a pipeline approach to predict important intermediate features (i.e. micro- and macro-level features) before predicting the overall quality score. This design also has another advantage: it is more interpretable because, given an overall score, we can also look at the intermediate results to understand what or where went wrong, providing more nuanced feedback that can benefit both teachers and second language students. Lastly, the strong performance of GPT4 (which is only marginally behind fine-tuned BERT) also has another important implication: we can adapt our SL assessment system to another language without requiring large-scale manually-annotated data by prompting LLMs.
Models F1
BERT (One-step) 0.379
GPT4 (One-step) 0.585
Human+LR+LR 0.772
Human+BERT+BERT 0.860
Human+GPT4+GPT4 0.842
BERT+LR+LR 0.329
BERT+BERT+BERT 0.807
GPT-4+LR+LR 0.667
GPT4+GPT4+GPT4 0.791
Table 7: F1 performance of the dialogue overall score for the CSL dialogue in three steps in different models
Models Topic Tone Opening Closing
LR 0.643 0.357 0.081 0.079
RF 0.463 0.235 0.066 0.090
NB 0.560 0.290 0.106 0.117
Table 8: Feature Importance of the four interactivity labels across three machine learning models
6.3 Additional Analyses
To understand how macro-level interactivity labels contribute to the overall quality score, we present their feature importance as learned by LR, RF and NB in the “Human+LR+LR”, “Human+RF+RF” and “Human+NB+NB” models in Table 8. Topic management appears to be the most important predictor, followed by tone appropriateness. Conversation opening and closing, on the other hand, has relatively small weights. These observations are supported by Roever and Ikeda (2023), who similarly found that topic flow and tone are dominant factors that decide dialogue quality.
Next, we assess how well BERT and GPT4 perform for the first (micro-level feature span prediction) and the second step predictions (macro-level label classification); full results in AppendixA.5 Table14 and Table15 for BERT and GPT4 respectively. To summarise, for most of the 17 micro-level features (step 1), both BERT and GPT4 perform very well (average F1 for BERT and GPT4 is 0.89 and 0.81, respectively). That said, BERT and GPT4 struggle on two features related to non-factive verbs, and this can be attributed to the rarity of these features in the Chinese Roever and Ikeda (2023). Overall, these results show that the proposed micro-features by Gao et al. (2024) transfer very well from English to Chinese.
For the second step, we look at the performance of “BERT+BERT” and “GPT4+GPT4” for predicting the macro-level labels (Appendix Table 16). Across all labels, both BERT and GPT4 perform well, averaging about 0.83 F1 for BERT and 0.78 F1 for GPT4, respectively. Together, these encouraging intermediate evaluation results explain why our pipeline approach can consistently score the overall dialogue quality.
7 Conclusion
We develop CNIMA, a CSL dataset with 10K dialogues annotated based on an evaluation framework that assesses micro-level features, macro-level interactivity labels, and overall quality scores. As the evaluation framework is originally designed for ESL dialogues Gao et al. (2024), our first contribution tests how well the framework transfers to CSL dialogues. We found that the evaluation framework is robust across languages and further revealed universal and language-specific insights about the relationships between micro-level and macro-level features. Our second contribution is we propose an automated approach that predicts the micro-level features, macro-level labels and overall quality scores in sequence. We experimented with classical models, BERT and GPT-4o, and found that the BERT and GPT4-o-based systems perform very well in predicting the overall dialogue quality score. Our approach is interpretable, and the LLM-based variant can be easily adapted to another language for second language assessment without requiring annotated training data.
8 Limitations
The original speech data is transcribed manually into text. Technically we can use speech-to-text software to automate this, but since we’re working with second language conversations, this can result in much noise in the translated text. Ultimately, this does mean fully automating second language assessment using interactive conversations remains a challenging task.
The scope of the current paper is limited to CSL, but it would be equally interesting to see how the evaluation framework would work for other SL dialogues. Admittedly the models we experiment with are limited in terms of novelty, but we contend that is not the focus of our contribution in this paper.
Ethics Statement
This study is conducted under the guidance of the ACL Code of Ethics. We manually filtered out potentially offensive content and removed all information related to the identification of annotators. The annotation protocol is approved under The University of Melbourne Human Ethics LNR as 2022-24988-32929-3.
Appendix A Appendix
A.1 Software Availability
To contribute to the research community and facilitate further development and collaboration, we have made the source codes of our innovative annotation tool publicly available. The tool, designed with a focus on enhancing the efficiency and accuracy of data annotation processes, has been developed through meticulous research and development efforts. It incorporates a range of features tailored to meet the needs of researchers and practitioners working in fields that require precise and reliable annotation of datasets.
Accessing the Source Code
The source codes are hosted on GitHub, a platform widely recognized for its robust version control and collaborative features. Interested parties can access the repository at the following link: https://anonymous.4open.science/r/AnnotationTool2023-CFE1/README.md. This repository is intended for research usage, underlining our commitment to supporting academic and scientific endeavours.
Key Features and Capabilities
Our annotation tool stands out for its user-friendly interface, which simplifies the annotation process and allows users to work more efficiently. Among its key features are:
  • Customizable Annotation Labels: Users can add their own set of labels to cater to the specific requirements of their projects.
  • Collaborative Annotation Support: Facilitating teamwork, the tool allows multiple annotators to work on the same dataset simultaneously, ensuring consistency and reducing the time required for project completion.
  • Annotation History Tracking: All the annotation history, such as changes made, can be tracked, and any further modifications can be done at any time for the user’s convenience. Export Functionality: Annotated data can be exported in several formats, accommodating further analysis or use in machine learning models.
A.2 Pages View For Span Annotation Website Interface
Figure 3: Annotation tool Demo
Figure 4: Hierarchical Label Assignment Demo
A.5 BERT Configuration
We give detailed experimental settings in the following table. All the experiments can run via a single NVIDIA GeForce RTX 4070 GPU within a reasonable time.
Table 9: Experimental Details
Parameters Configurations
Max Length 128
Tokenizer Bert Pretrained Tokenizer
Encoder BERT-base-uncase
Activation Function GELU
Batch Size 32
Learning Rate 5e-5
Loss Function CrossEntropyLoss
Optimizer Adam
Epoch 15
A.6 Prompts for GPT-4o
The following Table 10 shows the prompts for annotating micro-level features with GPT-4o. {CJK}UTF8gbsn
Span Description and Example
Reference Word A linguistic term used to avoid repetition and link different parts.
Example: [你]到哪里去 (Where are [you] going?)
Noun & Verb Collocation Words or phrases that habitually occur together, forming a strong and natural linguistic association.
Example:[帮我一个忙]吗? (Can you [do me a favor]?)
Code-switching The alternation between languages within a single conversation.
Example: 我觉得, [well], 这个是对的 (I think, well, this is right.)
Negotiation of Meaning Interactive process where speakers clarify and confirm understanding.
Example: SPK1: 那你是没有做吗?
SPK2: [的确], 我一点也没学 (SPK1: So you didn’t do it? SPK2: [Indeed], I haven’t learned at all.)
Tense Choice to Show Interactive Aims Using verb tenses to fulfil specific communicative goals.
Example:[现在]还不能, 因为我还有很多工作 (I can’t [right now] because I still have a lot of work.)
Routinized Resources Prefabricated linguistic elements are used to manage dialogue interactions efficiently.
Example: [你说的你] ([As you said])
Subordinate Clauses Clauses that provide additional information to the main clause.
Example: 你是数学好, [但是] 我是因为需要自己努力 (You are good at math, [but] because I need to work hard myself.)
Backchannels Brief responses from a listener, such as "uh-huh".
Example: [哎哎哎。] ([Yep, yep, yep.])
Question Responses Replies in a dialogue that directly or answer a preceding question.
Example: [SPK2 你既然有反驳的能力, 你还是自学吧]
(SPK2: [Since you have the self-learning skills, you shall teach yourself.])
Formulaic Responses Conventional phrases in dialogue to respond in familiar situations.
Example: [差不了多少]([Roughly the same])
Collaborative Finishes Instances in a dialogue where one speaker completes another speaker’s sentence or thought.
Example: SPK1:[好嘞再见啊] SPK2:[再见了您]
(SPK1: [Alright, See you.] SPK2:[Goodbye.])
Epistemic Copulas Phrases that express a speaker’s degree of certainty about a statement, often using verbs like "is" or "seems".
Example: [一个人去还是觉得有点变扭] ([I felt wired to go there alone])
Epistemic Modals Modal verbs or phrases that express a speaker’s judgment about the possibility, such as "might," "must,".
Example: [你应该自己学会这些中文知识的] ([You should learn these Chinese by yourself.])
Adjectives & adverbs of possibility Adjectives or adverbs to show possibility, like "Possibly".
Example: [我也许回去故乡] ([I maybe go back to hometown.])
Non-factive Verb Phrase Expressions that use verbs to convey statements without asserting them as true; verbs "think," "believe," or "seem."
Example: [我姑且能跟上吧] ([I can barely follow the progress.])
Impersonal Subject An impersonal subject (such as "it" or "there") is followed by a non-factive verb and a noun phrase, often express opinions.
Example: [这不好说吧] ([It’s hard to tell.])
Feedback in Next Turn Using next turn to respond other speaker.
Example: [我认为你有道理] ([You words make sense.])
Table 10: LLM Annotation Prompts for CSL Dialogue Span annotation for Mirco-level features
A.7 Prompts for GPT-4o Dialogue Overall Evaluation
The following Table 11 shows the prompts for dialogue overall quality score with GPT-4o.
Field Description
CONVERSATION A dialogue of second language Chinese conversation.
Output Fields score: The score of the interactivity of the Chinese second language dialogue (1 to 5). rationale: The reason why and how the score is made.
Evaluation Criteria 5: Smooth and fluent daily communication, easy and pleasant. 4: Somewhat less fluent communication, but the communication purpose is achieved. 3: Slightly awkward communication, such as not being able to immediately understand the other person’s question with hesitation. 2: Overall communication is not fluent and awkward, but some parts can be mutually understood. 1: Unable to accurately achieve the communication purpose, awkward conversation, failed to talk throughout the conversation.
Table 11: LLM Dialogue Overall Dialogue Quality Evaluation Prompts
A.8 Overall Score Description and Definitions
The following Table 12 shows the descriptions and definitions for dialogue’s overall quality score.
Scores Descriptions
5
Smooth and fluent daily communication,
easy and pleasant through the whole chat
4
Somewhat less fluent communication,
but the communication purpose is achieved
3
Slightly awkward communication in some places,
such as not being able to understand the other person’s question
2
Overall communication is not fluent and mostly awkward,
but some parts can be mutually understood
1
Unable to accurately achieve the communication purpose,
awkward conversation, and failed to talk throughout the conversation.
Table 12: Score description for overall dialogue quality
A.9 Four Interactivity Aspects Definitions and Descriptions
Table 13 shows the descriptions and definitions for dialogue’s overall quality score.
Interactivity
Macro-level Features
Definition
Topic Management
the strategies and techniques used
to control and navigate the flow of topics
Tone Choice Appropriateness
the suitability of the tone used in communication,
ensuring it aligns with the context, audience,
and purpose to convey the intended message
Conversation Opening
the initial interaction or exchange that begins a dialogue,
often setting the tone and context for the dialogue
Conversation Closing
the process of ending a dialogue or interaction,
which involves signaling the conclusion of the discussion,
summarizing key points, and often expressing a farewell
Table 13: Definitions of macro-level interactivity features, with higher score emphasising on natural, authentic interaction and active engagement in the dialogue
A.10 Top-k feature computation method
This is reproduced from Gao et al. (2024). Given that a trained LR, NB and RF classifier all provide weights to indicate the importance of each feature, for each classifier, we first compute common micro-level features $f_{\mathrm{c}}$ across the four interactivity labels:
[…]
where $\mathrm{topk}$ is a function that returns the best $k$ items given by their weights, $f_{\mathrm{topic}}$ denote the set of micro-level features with their weights for predicting the topic management interactivity label.
For micro-level features that are specific to each of the marco-level interactivity aspects. To that end, for each classifier we compute interactivity-specific features, e.g., for topic management, as follows:
[…]
Appendix B Experimental Result
Response Type Accuracy Precision Recall F1 Score
Epistemic Copulas 0.997 0.929 0.843 0.884
Formulaic Responses 0.976 0.875 0.781 0.825
Question-based Responses 0.986 0.834 0.591 0.892
Non-factive Verb Phrase Structure 0.999 0.000 0.000 0.000
Impersonal Subject + Non-factive Verb + NP 0.997 0.909 0.243 0.384
Reference Word 0.990 0.985 0.989 0.987
Routinized Resources 0.983 0.783 0.706 0.742
Noun & Verb Collocation in Proper Form 0.985 0.970 0.958 0.964
Collaborative Finishes 0.995 0.802 0.631 0.706
Tense Choice to Indicate Interactive Aims 0.994 0.957 0.930 0.943
Negotiation of Meaning 0.991 0.849 0.713 0.775
Code-switching for Communicative Purposes 0.999 0.976 0.954 0.965
Feedback in the Next Turn 0.972 0.842 0.833 0.838
Epistemic Modals 0.997 0.909 0.945 0.926
Backchannels 0.982 0.824 0.731 0.875
Subordinate Clauses 0.990 0.960 0.937 0.948
Adv & Adj Expressing 0.964 0.863 0.831 0.847
Table 14: Predicted performance of micro-level features on fine-tune BERT
Response Type Accuracy Precision Recall F1 Score
Epistemic Copulas 0.995 0.700 0.736 0.717
Formulaic Responses 0.951 0.559 0.579 0.669
Question-based Responses 0.972 0.500 0.426 0.460
Non-factive Verb Phrase Structure 0.999 0.000 0.000 0.000
Impersonal Subject + Non-factive Verb + NP 0.998 0.000 0.000 0.000
Reference Word 0.987 0.982 0.983 0.982
Routinized Resources 0.971 0.564 0.448 0.500
Noun & Verb Collocation in Proper Form 0.972 0.953 0.900 0.826
Collaborative Finishes 0.988 0.565 0.406 0.472
Tense Choice to Indicate Interactive Aims 0.990 0.915 0.826 0.768
Negotiation of Meaning 0.971 0.560 0.462 0.506
Code-switching for Communicative Purposes 0.999 1.000 0.941 0.869
Feedback in the Next Turn 0.951 0.679 0.741 0.708
Epistemic Modals 0.995 0.973 0.770 0.860
Backchannels 0.975 0.697 0.588 0.638
Subordinate Clauses 0.964 0.863 0.831 0.847
Adv & Adj Expressing 0.901 0.731 0.706 0.847
Table 15: F1 performance of micro-level span annotation by GPT-4o
Models Topic Tone Opening Closing
BERT (raw dialogue) 0.414 0.401 0.414 0.379
GPT-4 (raw dialogue) 0.553 0.533 0.585 0.557
BERT+BERT (on annotated data) 0.987 0.990 0.993 0.978
BERT+BERT (based on BERT predicted micro-level features) 0.836 0.855 0.836 0.830
GPT-4+GPT-4 (on annotated data) 0.761 0.749 0.812 0.809
Table 16: F1 performance of Marco-level four interactivity aspects’ score prediction across different model versions

### v2

@@
-CNIMA: A Universal Evaluation Framework and Automated Approach for Assessing Second Language Dialogues
+An Interpretable and Crosslingual Method for Evaluating Second-Language Dialogues
 Abstract
-We develop CNIMA (Chinese Non-Native Interactivity Measurement and Automation), a Chinese-as-a-second-language labelled dataset with 10K dialogues.
-We annotate CNIMA using an evaluation framework — originally introduced for English-as-a-second-language dialogues — that assesses micro-level features (e.g.
+We analyse the cross-lingual transferability of a dialogue evaluation framework that assesses the relationships between micro-level linguistic features (e.g.
 backchannels) and macro-level interactivity labels (e.g.
-topic management) and test the framework’s transferability from English to Chinese.
-We found the framework robust across languages and revealed universal and language-specific relationships between micro-level and macro-level features.
-Next, we propose an approach to automate the evaluation and find strong performance, creating a new tool for automated second language assessment.
-Our system can be adapted to other languages easily as it uses large language models and as such does not require large-scale annotated training data.
-CNIMA: A Universal Evaluation Framework and
-Automated Approach for Assessing Second Language Dialogues
+topic management), originally designed for English-as-a-second-language dialogues.
+To this end, we develop CNIMA (Chinese Non-Native Interactivity Measurement and Automation), a Chinese-as-a-second-language labelled dataset with 10K dialogues.
+We found the evaluation framework to be robust across distinct languages: English and Chinese, revealing language-specific and language-universal relationships between micro-level and macro-level features.
+Next, we propose an automated, interpretable approach with low data requirement that scores the overall quality of a second-language dialogue based on the framework.
+Our approach is interpretable in that it reveals the key linguistic and interactivity features that contributed to the overall quality score.
+As our approach does not require labelled data, it can also be adapted to other languages for second-language dialogue evaluation.
+An Interpretable and Crosslingual Method for Evaluating Second-Language Dialogues
 Rena Gao, Jingxuan Wu, Carsten Roever, Xuetong Wu, Jing Wu, Long Lv, Jey Han Lau The University of Melbourne, Parkville, 3052, Australia The Chinese University of Hong Kong, Shatin, NT, Hong Kong, China Central South University, Changsha, 410017, China Tiangong University, Tianjin, 300387, China {rena.gao,carsten}@unimelb.edu.au, {wfyitf,jeyhan.lau}@gmail.com wujingxuan@cuhk.edu.cn, wujing@tiangong.edu.cn, lvlong@csu.edu.en
 1 Introduction
-In the context of second language (SL) assessment, speaking has always been considered an essential ability for SL speakers Smith et al.
+In the context of second language (SL) assessment, speaking has always been considered an essential ability Smith et al.
 (2022); Allwood (2008); Huang et al.
-(2020), but prior studies have predominantly focused on written language proficiency Paiva et al.
+(2020), but prior studies have predominantly focused on written correction Paiva et al.
+(2022) or pronunciation from ASR McGhee et al.
 (2024).
-Given the lack of datasets that capture the unique linguistic features of SL speakers in dialogues (especially in open-domain conversations), this leaves conversational interaction assessment under-explored and ultimately contribute to a limited understanding of conversational fluency and interactivity for SL speakers.
+Given the lack of researches that capture the unique linguistic features of SL speakers in dialogues (especially in open-domain interactive conversations), this leaves conversational interaction assessment under-explored and ultimately contributes to a limited understanding of conversational fluency and interactivity for SL speakers from different languages.
 An exception is Gao et al.
 (2025), who propose a two-level framework for assessing the interactivity ability of English-as-a-second-language (ESL) speakers in open-domain conversations.
-They introduce micro-level word/utterance features (e.g.
+The framework introduces micro-level word/utterance features (e.g.
 backchannels) and macro-level interactivity labels (e.g.
-topic management) and found that the micro-level features are highly predictive of the macro-level labels.
-Gao et al.
-(2024) reveal their intricate relationships, showing that certain micro-level features are highly correlated with a particular macro-level feature, e.g.
-frequent use of appropriate collocations and varied sentence structures (micro-level features) significantly correlate with topic management and tone appropriateness (macro-level labels).
-However, this research is conducted only in the context of ESL, raising questions about the transferability of the evaluation framework to other languages besides English.
-Furthermore, they did not propose an automated pipeline for assessing SL dialogue quality, as their predictive models rely on human-annotated micro-level features.
-Our work aims to address these shortcomings by (1) developing and testing a Chinese-as-a-second-language (CSL) dialogue dataset using the evaluation framework of Gao et al.
-(2024); (2) introducing a fully automated pipeline to assess interactivity for CSL conversations.
-We experiment with different machine learning models, from classical machine learning models (e.g.
-logistic regression) to large language models (LLMs; e.g.
-GPT-4o OpenAI (2024)), to automate the prediction of micro/macro-level features and overall dialogue quality scores.
-Our best system shows strong predictive performance, paving the way for a new tool for assessing SL conversations [^1] The dataset and related codes can be accessed at https://github.com/RenaGao/CSL2024.
-To summarise:
-  • We release CNIMA (Chinese Non-Native Interactivity Measurement and Automation), an annotated CSL dialogue dataset with 10K dialogues, based on the framework of Gao et al.
-(2024) originally designed for ESL.
-  • We test the cross-lingual transferability of the evaluation framework for CSL and find that it is robust across languages, and we further reveal universal and language-specific relationships between the micro-level and macro-level features.
-  • We introduce a fully automated approach that predicts micro-level features, macro-level labels and an overall quality score given an input dialogue by exploring a spectrum of models from classical machine learning to LLMs.
-Our best system demonstrates strong performance, creating a new automatic tool for second language assessment.
-Importantly, our LLM-based system can be adapted to other languages without requiring annotated data.
+topic management) and they find that the micro-level features are highly predictive of the macro-level labels.
+However, they conduct the study only in the context of ESL, raising questions about the transferability of the evaluation framework to other languages besides English.
+Furthermore, they do not introduce an automated approach for this evaluation, as the analysis relies on human annotations.
+Our work aims to address these shortcomings by: (1) testing the cross-lingual transferability of the evaluation framework by validating it on an annotated Chinese-as-a-second-language (CSL) dialogue dataset; and (2) introducing an automated, interpretable approach with low data requirement that scores the overall quality of a second-language dialogue.
+Our automated approach is interpretable as it highlights the key features that lead to the overall score, and it has a low data requirement because it does not require labelled data.
+As it can be adapted to other languages easily, our method paves the way for a new tool for assessing SL conversations.
+[^1] The dataset and related codes can be accessed at https://github.com/RenaGao/CSL2024 To summarise our contributions:
+  • We release CNIMA (Chinese Non-Native Interactivity Measurement and Automation), an annotated CSL dialogue dataset with 10K dialogues, based on the evaluation framework of Gao et al.
+(2025) originally designed for ESL dialogues.
+  • We evaluate the cross-lingual transferability of the evaluation framework and find it to be robust across different languages.
+We further reveal language-specific and language-universal relationships between micro-level and macro-level features, highlighting the subtle differences between CSL and ESL dialogues.
+  • We introduce an automated, interpretable approach that predicts the micro- and macro-level features and the overall quality scores of second-language dialogues.
+Our best method demonstrates strong performance, creating a new tool for second language assessment.
+Importantly, the method can be adapted to other languages as it does not require labelled data.
 2 Related work
@@
 (2021).
-However, few speaking assessments emphasised the importance of interaction Khabbazbashi et al.
-(2021) and aspects such as how speakers manage the topics in communication Shaxobiddin (2024), how speakers perform social roles from speaking Chen et al.
+However, few speaking assessments emphasised the importance of interaction in dialogues for second language speakers and learners Khabbazbashi et al.
+(2021) and aspects such as how speakers manage the topics in communication Shaxobiddin (2024), how speakers perform social roles from speaking interactions Chen et al.
 (2023), and how speakers start and end a talk in an acceptable manner Yap and Sahoo (2024).
 There are, however, some exceptions.
-For example, Dai (2022) develop a test rubric for Chinese SL speakers.
+For example, Dai (2022) develop a test rubric for Chinese Second Language speakers.
 More recently, Gao and Wang (2024) introduce an interactivity scoring framework inspired by the IELTS speaking assessment.
-However, these studies only provide a theoretical framework for evaluation and lack an automated pipeline for large-scale assessments.
+However, these studies only provide a theoretical framework for evaluation and lack an automated pipeline for large-scale assessments, human rating is still a major scoring approach in mainstream industrial practice.
 2.2 Automated Scoring System in SL
 Automated scoring systems can improve the efficiency of processing and scoring dialogues compared to manual assessment methods, allowing for real-time feedback Evanini et al.
-(2017).
+(2017) and explainable suggestions.
 Some major second language assessment organisations have employed automated speaking scoring, including PTE Pearson Jones and Liu (2023), Duolingo English Test Burstein et al.
-(2021), and TOEFL Gong (2023), but these automated systems fail to capture the interactive nature of SL conversations.
+(2021), and TOEFL Gong (2023), but these automated systems failed to capture the interactive nature of SL conversations.
 That is, they struggle to offer detailed analyses and insights on the common errors and language usage patterns and where SL speakers struggle the most.
-This motivates a more explainable framework for SL assessment that can provide better feedback.
-The main challenge is in designing an effective evaluation framework.
-Capturing the nuances of spoken language, such as interaction, attitudes, and cooperation in conversation, is complex to achieve Gao and Wang (2024).
-Automated scoring systems must handle diverse patterns, which can vary widely among different non-native SL speakers.
-Recognizing and assessing interactive features like turn-taking, interruptions, and response appropriateness further complicates the process.
-Ultimately, the dynamic nature of conversations makes it challenging for automated systems to accurately evaluate interactive speaking, necessitating sophisticated algorithms and continuous refinement Cumbal (2024); Engwall et al.
-(2022).
+This motivates a more explainable framework for SL assessment that can provide better feedback and suggestions from the automated scoring.
+The main challenge is to design an effective automated evaluation tool.
+Capturing the nuances of spoken language, such as interaction, attitudes, and cooperation in conversation, is challenging to achieve Gao and Wang (2024).
+Automated scoring systems must handle diverse patterns across varied languages, which can vary widely among different non-native SL speakers across distinct languages, like Asian languages and European languages with distinct differences.
+Assessing interactive features like turn-taking, interruptions, and response appropriateness further complicates the process.
+Ultimately, the dynamic nature of conversations makes it challenging for automated systems to accurately evaluate interactive speaking, necessitating sophisticated algorithms and continuous refinement Engwall et al.
+(2022); Cumbal (2024).
 3 Evaluation Framework
-We adapt the dialogue evaluation framework of Gao et al.
-Gao et al.
-(2024) for assessing SL dialogues, which considers micro-level features and macro-level interactivity labels.
+We extend the evaluation framework of Gao et al.
+(2025) for assessing second-language dialogues, which considers micro-level features and macro-level interactivity labels.
 One addition we have made is to introduce an overall dialogue quality score.
@@
 A more elaborate definition can be found in Appendix Table 13.
-For each macro-level interactivity label, the score ranges from 1 to 5 (categorical), and higher scores indicate a more natural and authentic interactivity quality.
+For each macro-level interactivity label, the score ranges from 1 to 5 (categorical), and higher scores indicate a more natural and active interactivity quality.
 Note that for tone choice appropriateness, higher scores denote a casual tone, and lower scores indicate a formal tone.
@@
 We followed a similar process, where we recruited 20 learners of Chinese Mandarin at four different proficiency levels.
-[^3] Proficiency level can be categorized into the following levels: beginner, lower, intermediate, and high Wu and Roever (2021).
+[^3] Proficiency level can be categorized into beginner, lower, intermediate, and high Wu and Roever (2021).
 Each participant was assigned to a dyadic group with a partner of a similar proficiency level.
@@
 [^5] The workers are undergraduate linguistics major students, and they are native Chinese speakers.
-Including the original 8,000 dialogues from Wu and Roever (2021), our extended CSL dataset has 10,908 dialogues in total; table 1 presents some statistics.
+Including the original 8,000 dialogues from Wu and Roever (2021), our extended CSL dataset has 10,908 dialogues in total as shown in Table 1.
 Statistic Count
@@
 Table 2: Inter-annotator agreement for micro-level, macro-level features and overall scores.
-4.2 Macro/Micro-level Feature Annotation
+4.2 Data Annotation
 Given the dialogues, we annotate them for micro-level features, macro-level interactivity labels and overall quality scores based on the evaluation framework introduced in Section 3.
-To this end, we have recruited twelve postgraduate students who are native Chinese speakers.
-The authors of this paper first trained the annotators and the annotator training manual can be found in Appendix  LABEL:subapp:MN.
+To this end, we recruited twelve postgraduate students who are native Chinese speakers.
+We (first author) first trained the annotators and the annotator training manual can be found in Appendix  LABEL:subapp:MN.
 During the training, annotators will see one dialogue as an example to understand the requirements and learn how to use the annotation platform (Appendix A.1).
@@
 For the overall score, in addition to providing the score, the annotators are also asked to write down their reasons for justifying their label.
-As explained in Section 3, the micro-level features are annotated as spans, and the macro-level labels and overall quality scores are document (dialogue) labels.
+As explained in Section 3, the micro-level features are annotated as spans, and the macro-level labels and overall quality scores as document (dialogue) labels.
 To aggregate the span annotations by the two annotators for the micro-level features, for each dialogue and each micro-level feature, we iterate through each turn and select the shorter (longer) span between the two annotators if it is a token-level (utterance-level) feature.
@@
 The majority label is the most frequent label when we pool together all the labels from the segmented, smaller dialogues that belong to the same original conversation.
-Our rationale for deferring to this majority label is that these are cases where it is difficult to determine the correct dialogue level label using the smaller segmented dialogue (this happens more with topic management than other interactivity labels), and so it is best that we look at all the labels across all the segmented dialogues collection that belong to the same conversation.
+Our rationale for deferring to this majority label is that these are cases where it is difficult to determine the correct dialogue level label using the smaller segmented dialogue (this happens more with topic management than other interactivity labels), and so we look at all the labels across all the segmented dialogues collection that belong to the same conversation.
 For the overall dialogue quality scores, for each dialogue with disagreement, we (two authors of this paper) manually assess the justification provided by the annotators to determine the ground truth.
@@
 This is so we create an unbiased ground truth for the overall score that is independent of the macro-level labels.
-To measure annotation quality, for the micro-level features, we measure agreement between the annotators at the token level for each micro-level feature, i.e., we compute agreement statistics based on the presence or absence of the feature as marked by the annotators for each word token.
+To measure the annotation quality, for the micro-level features, we calculate agreement between the annotators at the token level for each micro-level feature, i.e., we compute agreement statistics based on the presence or absence of the feature as marked by the annotators for each word token.
 [^10] In other words, the unit of analysis here is a word token, and the output is a binary value for each annotator indicating whether it has been marked for the feature.
 We calculate Pearson correlation coefficient $r$ Cohen et al.
-(2009) and Krippendorff’s $\alpha$ Krippendorff (2018) summarise the results in Table 2.
-The agreement is above 0.6 over all the features/labels, indicating a good consensus between annotators under this framework.
+(2009) and Krippendorff’s $\alpha$ Krippendorff (2018) and summarise the results in Table 2.
+The agreement is above 0.6 over all the features/labels, indicating a good consensus between annotators.
 Figure 2: Pipeline for automated scoring of the CSL dialogue on three steps
-5 Automation Pipeline
+5 Automated Pipeline
 We now propose automating the prediction of the micro-level features (step 1), macro-level interactivity labels (step 2), and overall score (step 3); see Figure 2 for an illustration.
@@
 This step predicts the macro-level interactivity labels for topic management, tone choice appropriateness, and conversation opening and closing.
-Here, we experiment with classical models (LR, RF and NB), BERT and GPT-4o, and for each of them, we train four (one for each interactivity label) 5-class classifiers (each interactivity label has 5 classes).
+Here, we experiment with classical models (LR, RF and NB), BERT and GPT-4o.
+For each of them, we train four (one for each interactivity label) 5-class classifiers (each label has 5 classes).
 For the classical models, we follow Gao et al.
@@
 This step predicts the overall dialogue quality score.
-As this is a document classification like step 2, we follow largely the same process for building LR, RF, NB, BERT and GPT-4o.
+As this is a document classification task like step 2, we follow the same process for building LR, RF, NB, BERT and GPT-4o.
 Note that the classical models (LR, RF, NB) use only the normalised macro-level interactivity score as input features (in other words, they have only 4 input features), while BERT and GPT-4o uses a concatenation of the dialogue, micro-level feature spans, and macro-level interactivity scores as input.
@@
 We first replicate the ESL experiments of Gao et al.
-(2024) for CSL and compare their findings in Section 6.1.
+(2025) for CSL and compare their results in Section 6.1.
 We then assess the performance of our proposed automated approach for SL assessment in Section 6.2.
@@
 6.1 Transferability of the Evaluation Framework from ESL to CSL
-Based on the human-annotated micro-level features, Gao et al.
-(2024) convert them into normalised counts and use them as input features to train an LR, NB and RF classifier to predict macro-level interactivity labels (i.e.
-topic management, tone appropriateness, conversation opening and closing).
+For the human-annotated micro-level features, Gao et al.
+(2025) convert them into normalised counts and use them as input features to train an LR, NB and RF classifiers to predict macro-level interactivity labels.
 They found: (1) strong prediction performance; (2) high-impact micro-level features that are common across all interactivity labels (by interpreting feature importance given by the trained classifiers); and (3) high-impact micro-level features that are specifically predictive for an interactivity label.
-We follow Gao et al.
-(2024) to compute the common and specific features for (2) and (3); details in Appendix A.10.
-We replicate their experiments here, using our annotated CSL data (CNIMA), to see whether their findings transfer across languages.
+In the following, we replicate the experiments using our annotated CSL data (CNIMA), to see whether the results transfer across languages; details can be found in Appendix A.10.
 For (1), we present the F1 performance of predicting macro-level interactivity labels for CNIMA in Table 3.
@@
 6.2 Evaluation of Automated Pipeline
-We now evaluate our automated 3-step approach to predicting the overall dialogue quality score and present the results in Table 7.
+We now evaluate our automated 3-step approach to predicting the overall dialogue quality score as results in Table 7.
+Our pipeline predicts micro- and macro-level features and dialogue overall quality scores.
 In terms of model names, each component denotes the model used in a particular step, e.g.
@@
 Overall, the performance is encouraging, and we see that the best models achieve over 0.80 F1.
-When we look at the fully automated pipeline (bottom 4 rows in Table 7), “‘BERT+BERT+BERT” and “‘GPT4+GPT4+GPT4” perform very strongly (0.807 and 0.791), demonstrating that we have a fully automated system that can reliably assess the quality score of a dialogue.
+When we look at the fully automated pipeline (bottom 4 rows in Table 7), “BERT+BERT+BERT” and “GPT4+GPT4+GPT4” perform very strongly (0.807 and 0.791), demonstrating that we have a fully automated system that can reliably assess the quality score of a dialogue.
 We also notice an interesting trend: When we use either BERT or GPT4 for predicting the micro-level features, LR (i.e.
-“BERT+LR+LR” and “GPT4+LR+LR”) performs very poorly for predicting the overall score, even though the span prediction performance of BERT and GPT4 for the micro-level features is no+t poor (we will revisit this in Section 6.3).
+“BERT+LR+LR” and “GPT4+LR+LR”) performs very poorly for predicting the overall score, even though the span prediction performance of BERT and GPT4 for the micro-level features is not poor (we will revisit this in Section 6.3).
 This suggests that the classical machine learning models are less tolerant of noise.
@@
 Taking all these results together, we show that when it comes to assessing SL dialogue quality, it is important to take a pipeline approach to predict important intermediate features (i.e.
-micro- and macro-level features) before predicting the overall quality score.
-This design also has another advantage: it is more interpretable because, given an overall score, we can also look at the intermediate results to understand what or where went wrong, providing more nuanced feedback that can benefit both teachers and second language students.
-Lastly, the strong performance of GPT4 (which is only marginally behind fine-tuned BERT) also has another important implication: we can adapt our SL assessment system to another language without requiring large-scale manually-annotated data by prompting LLMs.
+micro- and macro-level) before predicting the overall quality score.
+This design also has another advantage: it is more interpretable as given an overall score, we can also look at the intermediate results to understand what and where went wrong, providing more meaningful feedback that can benefit both teachers and second language students.
+Lastly, the strong performance of GPT4 also has another important implication: we can adapt our SL assessment system to another language without requiring large-scale manually-annotated data.
 Models F1
@@
 BERT+BERT+BERT 0.807
-GPT-4+LR+LR 0.667
+GPT4+LR+LR 0.667
 GPT4+GPT4+GPT4 0.791
@@
 6.3 Additional Analyses
-To understand how macro-level interactivity labels contribute to the overall quality score, we present their feature importance as learned by LR, RF and NB in the “Human+LR+LR”, “Human+RF+RF” and “Human+NB+NB” models in Table 8.
+To understand how macro-level interactivity labels contribute to the overall quality score, we present their feature importance as learned by classical machine learning models in the “Human+LR+LR”, “Human+RF+RF” and “Human+NB+NB” models in Table 8.
 Topic management appears to be the most important predictor, followed by tone appropriateness.
 Conversation opening and closing, on the other hand, has relatively small weights.
-These observations are supported by Roever and Ikeda (2023), who similarly found that topic flow and tone are dominant factors that decide dialogue quality.
-Next, we assess how well BERT and GPT4 perform for the first (micro-level feature span prediction) and the second step predictions (macro-level label classification); full results in AppendixA.5 Table14 and Table15 for BERT and GPT4 respectively.
+These observations are supported by Roever and Ikeda (2023), who similarly found that topic flow and tone are dominant in deciding interaction quality.
+Next, we assess how well BERT and GPT4 perform for the first (micro-level feature span prediction) and the second-step predictions (macro-level label classification); full results in AppendixA.5 Table14 and 15 for BERT and GPT4 respectively.
 To summarise, for most of the 17 micro-level features (step 1), both BERT and GPT4 perform very well (average F1 for BERT and GPT4 is 0.89 and 0.81, respectively).
 That said, BERT and GPT4 struggle on two features related to non-factive verbs, and this can be attributed to the rarity of these features in the Chinese Roever and Ikeda (2023).
-Overall, these results show that the proposed micro-features by Gao et al.
-(2024) transfer very well from English to Chinese.
+Overall, these results show that the proposed micro-features transfer very well from English to Chinese.
 For the second step, we look at the performance of “BERT+BERT” and “GPT4+GPT4” for predicting the macro-level labels (Appendix Table 16).
@@
 7 Conclusion
-We develop CNIMA, a CSL dataset with 10K dialogues annotated based on an evaluation framework that assesses micro-level features, macro-level interactivity labels, and overall quality scores.
-As the evaluation framework is originally designed for ESL dialogues Gao et al.
-(2024), our first contribution tests how well the framework transfers to CSL dialogues.
-We found that the evaluation framework is robust across languages and further revealed universal and language-specific insights about the relationships between micro-level and macro-level features.
-Our second contribution is we propose an automated approach that predicts the micro-level features, macro-level labels and overall quality scores in sequence.
-We experimented with classical models, BERT and GPT-4o, and found that the BERT and GPT4-o-based systems perform very well in predicting the overall dialogue quality score.
-Our approach is interpretable, and the LLM-based variant can be easily adapted to another language for second language assessment without requiring annotated training data.
+We introduce CNIMA, a large-scale annotated dataset for Chinese-as-a-Second-Language (CSL) dialogues based on the evaluation framework introduced by Gao et al.
+(2025).
+This allows us to test the cross-lingual transferability of the framework, as it is originally designed for English.
+Our experiments demonstrate that the framework works well for CSL dialogues, suggesting that it is robust across languages and that it is a viable framework for evaluating any second-language conversations.
+Our analyses also reveal language-specific and language-universal relationships between micro-level linguistic features (such as backchannels) and macro-level interactivity features (such as topic management).
+Next, we propose an automated, interpretable approach that predicts the micro- and macro-level features and the overall quality of second-language dialogues.
+To this end, we explore a range of models, from classical machine learning models to LLMs.
+Our approach is interpretable because it does not only predict the overall dialogue quality score, but also reveals the key micro-level and macro-level features that contribute to the score.
+Our LLM-based approach performs very well, and as it does not require any labelled data for training, it can be adapted to other languages easily.
+Ultimately, our method paves the way for an interpretable and practical system for automatic second-language dialogue assessment.
 8 Limitations
 The original speech data is transcribed manually into text.
-Technically we can use speech-to-text software to automate this, but since we’re working with second language conversations, this can result in much noise in the translated text.
-Ultimately, this does mean fully automating second language assessment using interactive conversations remains a challenging task.
-The scope of the current paper is limited to CSL, but it would be equally interesting to see how the evaluation framework would work for other SL dialogues.
-Admittedly the models we experiment with are limited in terms of novelty, but we contend that is not the focus of our contribution in this paper.
+Technically we can use speech-to-text software to automate this, but since we are working with second language conversations, this can result in much noise in the translated text.
+The scope of the current paper is limited to CSL, but it would be equally interesting to see how the evaluation framework would work for other SL dialogues, which calls on a joint contribution of the community effort on more SL language dialogue datasets, like Japanese second language datasets, Spanish second language datasets and etc.
 Ethics Statement
-This study is conducted under the guidance of the ACL Code of Ethics.
-We manually filtered out potentially offensive content and removed all information related to the identification of annotators.
-The annotation protocol is approved under The University of Melbourne Human Ethics LNR as 2022-24988-32929-3.
+This study is conducted under the guidance of the ACL Code of Ethics. We manually filtered out potentially offensive content and removed all information related to the identification of annotators.
 Appendix A Appendix
@@
 Negotiation of Meaning Interactive process where speakers clarify and confirm understanding.
-Example: SPK1: 那你是没有做吗?
-SPK2: [的确], 我一点也没学 (SPK1: So you didn’t do it? SPK2: [Indeed], I haven’t learned at all.)
+Example: SPK1: 那你是没有做吗? SPK2: [的确], 我一点也没学 (SPK1: So you didn’t do it? SPK2: [Indeed], I haven’t learned at all.)
 Tense Choice to Show Interactive Aims Using verb tenses to fulfil specific communicative goals.
@@
 Question Responses Replies in a dialogue that directly or answer a preceding question.
-Example: [SPK2 你既然有反驳的能力, 你还是自学吧]
-(SPK2: [Since you have the self-learning skills, you shall teach yourself.])
+Example: [我还是自学吧] (SPK2: [I shall teach myself.])
 Formulaic Responses Conventional phrases in dialogue to respond in familiar situations.
@@
 Conversation A dialogue of second language Chinese conversation.
-Output Fields score: The score of the interactivity of the Chinese second language dialogue (1 to 5). rationale: The reason why and how the score is made.
+Output Fields score: The score of the interactivity of the Chinese second language dialogue (1 to 5).
+rationale: The reason why and how the score is made.
 Evaluation Criteria 5: Smooth and fluent daily communication, easy and pleasant.